# Towards Widening The Distillation Bottleneck for Reasoning Models

### v1

@@ Abstract
Large Reasoning Models (LRMs) such as OpenAI o1 and DeepSeek-R1 have shown remarkable reasoning capabilities by scaling test-time compute and generating long Chain-of-Thought (CoT). Distillation—post-training on LRMs-generated data—is a straightforward yet effective method to enhance the reasoning abilities of smaller models, but faces a critical bottleneck: we found that distilled long CoT data poses learning difficulty for small models and leads to the inheritance of biases (i.e. over-thinking) when using Supervised Fine-tuning (SFT) and Reinforcement Learning (RL) methods. To alleviate this bottleneck, we propose constructing tree-based CoT data from scratch via Monte Carlo Tree Search (MCTS). We then exploit a set of CoT-aware approaches, including Thoughts Length Balance, Fine-grained DPO, and Joint Post-training Objective, to enhance SFT and RL on the construted data. We conduct evaluation on various benchmarks such as math (GSM8K, MATH, AIME). instruction-following (Multi-IF) and planning (Blocksworld), results demonstrate our approaches substantially improve the reasoning performance of distilled models compared to standard distilled models via reducing the hallucinations in long-time thinking.
Towards Widening The Distillation Bottleneck for Reasoning Models
Huifeng Yin1, 2thanks: Equal contribution.  Yu Zhao [^1] footnotemark: 1  Minghao Wu1, 3  Xuanfan Ni1  Bo Zeng1  Hao Wang1  Tianqi Shi1 Liangying Shao1  Chenyang Lyu1  Longyue Wang1thanks: Email: wanglongyue.wly@alibaba-inc.com.  Weihua Luo1  Kaifu Zhang1 1Alibaba International Digital Commerce 2Tsinghua University 3Monash University
@@ 2.4 CoT Data Extraction for Post-Training
  • CoT Data for SFT We typically select successful paths that lead to the correct final answer. Depending on the data volume requirements, one can: 1) Pick the highest-reward path according to MCTS; 2) Pick the long or short path that yields the correct answer, if specific chain lengths are desired.
  • CoT Data for DPO Constructing DPO data requires both positive and negative examples for each prompt: 1) The positive example is the CoT path that correctly solves the problem, like the SFT data; 2) The negative example is a flawed path (an incorrect final answer) that shares a minimal prefix with the positive path to mitigate excessive overlapping tokens, which can degrade DPO performance.
@@ 3.1 Thoughts Length Balance
Section 2.4 shows that the length of CoT significantly affects reasoning performance of distilled smaller models at DPO phase. However, in preliminary experiments, we found no such effect on SFT training. Therefore, we propose using longest CoT data during the SFT phase, while using the shortest CoT data during the DPO phase. In practice, we extract all valid paths leading to correct answer nodes from the CoT trees, as there can be multiple correct paths. From these paths, we select examples of varying lengths (long, medium, short) based on token count for SFT. For DPO, the correct paths serve as positive examples, while negative examples are generated by identifying incorrect paths that share the shortest common prefix with the positive examples. This ensures a diverse set of reasoning paths for both SFT and DPO.
@@ 4.2 Experimental Results
We apply our constructed CoT data to smaller models such as Llama-8B, comparing it against the Sky-T1 8B model, which utilizes a distillation pipeline based on QwQ and proves effective on larger models (32B). While Sky-T1 demonstrates competitive performance on large models, it faces challenges when scaled down to 8B models due to the inherent limitations in context processing and reasoning capabilities. In contrast, our CoT data, specifically designed to address these limitations, leads to substantial improvements in smaller models, particularly in tasks involving arithmetic reasoning such as GSM8K and MATH, as well as more complex open-ended tasks like AIME and Blocksworld, as shown in Table 4. This validates the effectiveness of our constructed data in advancing the performance of small models across a wide range of reasoning tasks.
@@ 4.4 MCTS Inference Exploration
As an additional experiment, we explored the impact of applying MCTS at the inference stage. As shown in Table 7, we use Test@N to denote the percentage of problems solved correctly at least once when allowing the model to make N separate guesses for each problem.Cobbe et al. (2021b) We evaluated solve rates at Test@1, Test@8, and Test@32 on the MATH dataset. Specifically, at Test@8 and Test@32, the MCTS-based approach outperforms the model without MCTS inference, demonstrating its ability to expand the solution space and leverage test-time scaling effectively.
@@ 5 Related Work
Researchers focus on transferring the reasoning capabilities of LRMs into smaller models, often through distillation techniques. Distillation methods, such as fine-tuning smaller models on CoT data generated by larger models Huang et al. (2024a), show that small models can benefit from reasoning data generated by large models. However, recent findings reveal that small models often fail to capture intricate reasoning patterns due to their limited capacity, leading to suboptimal performance when directly distilled from large models Li et al. (2025); Wang et al. (2025a).
@@ 6 Conclusion
We explore strategies for transferring long CoT reasoning to smaller models, addressing learning difficulties and bias inheritance common in standard distillation. We propose a MCTS framework that generates tree-based reasoning paths from scratch, enabling flexible data generation and reducing reliance on large teacher models. We enhance performance through CoT-aware post-training, combining supervised fine-tuning, direct preference optimization, and masking-based strategies. Experiments on diverse benchmarks-covering exam tasks, multilingual instruction following, and real-world planning-show that MCTS-generated data and targeted optimization reduce formalistic overthinking and content repetition. By balancing reasoning length, introducing multi-model coordination for reflection, and using a joint loss approach for preference optimization, our model demonstrates robust performance across various reasoning tasks. These findings highlight the importance of well-designed data and post-training strategies in improving the efficiency and reliability of smaller-scale reasoning models.
Limitations
Our MCTS-based method for constructing CoT data moves beyond straightforward distillation pipelines by explicitly building a search tree, guiding the LLM to generate multi-step, human-like reasoning. Despite its advantages, the method can be computationally expensive as task complexity increases. Future optimizations could include: 1) Pruning Strategies: Prune branches of the tree that appear suboptimal based on a partially learned reward or confidence metric. 2) More Diverse Model Rotations: Beyond reflection vs. thinking, multiple specialized models (or specialized parameter shards within a large model) could be employed for different reasoning skills (e.g., logic vs. creative).
Ethics Statement
We take ethical considerations very seriously, and strictly adhere to the ACL Ethics Policy. All datasets used in this paper are publicly available. We ensure that the findings and conclusions of this paper are reported accurately and objectively.

### v2

# Marco-o1 v2: Towards Widening The Distillation Bottleneck for Reasoning Models

@@ Abstract
Large Reasoning Models (LRMs) such as OpenAI o1 and DeepSeek-R1 have shown remarkable reasoning capabilities by scaling test-time compute and generating long Chain-of-Thought (CoT). Distillation post-training on LRMs-generated data is a straightforward yet effective method to enhance the reasoning abilities of smaller models, but faces a critical bottleneck: we found that distilled long CoT data poses learning difficulty for small models and leads to the inheritance of biases (i.e. over-thinking) when using Supervised Fine-tuning (SFT) and Reinforcement Learning (RL) methods. To alleviate this bottleneck, we propose constructing tree-based CoT data from scratch via Monte Carlo Tree Search (MCTS). We then exploit a set of CoT-aware approaches, including Thoughts Length Balance, Fine-grained DPO, and Joint Post-training Objective, to enhance SFT and RL on the constructed data. We conduct evaluation on various benchmarks such as math (GSM8K, MATH, AIME). instruction-following (Multi-IF) and planning (Blocksworld), results demonstrate our approaches substantially improve the reasoning performance of distilled models compared to standard distilled models via reducing the hallucinations in long-time thinking. The project homepage is https://github.com/AIDC-AI/Marco-o1.
Marco-o1 v2: Towards Widening The Distillation Bottleneck for Reasoning Models
Huifeng Yin1, 2thanks: Equal contribution.  Yu Zhao [^1] footnotemark: 1  Minghao Wu1, 3  Xuanfan Ni1  Bo Zeng1  Hao Wang1  Tianqi Shi1 Liangying Shao1  Chenyang Lyu1  Longyue Wang1thanks: Corresponding author.  Weihua Luo122footnotemark: 2  Kaifu Zhang1 1Alibaba International Digital Commerce 2Tsinghua University          3Monash University {fengli.zy, wanglongyue.wly, weihua.luowh, kaifu.zkf}@alibaba-inc.com
@@ 2.4 CoT Data Extraction for Post-Training
  • CoT Data for SFT. We select successful paths that lead to the correct final answer. Depending on the data volume requirements, one can: 1) Pick the highest-reward path according to MCTS; 2) Pick the long or short path that yields the correct answer, if specific chain lengths are desired.
  • CoT Data for DPO. Constructing DPO data requires both positive and negative examples for each prompt: 1) The positive example is the CoT path that correctly solves the problem, like the SFT data; 2) The negative example is a flawed path (an incorrect final answer) that shares a minimal prefix with the positive path to mitigate excessive overlapping tokens, which can degrade DPO performance.
@@ 3.1 Thoughts Length Balance
Section 2.4 illustrates that the length of Chain-of-Thought (CoT) reasoning significantly influences the reasoning performance of distilled smaller models during the Direct Preference Optimization (DPO) phase. In contrast, our preliminary experiments did not indicate such an effect during SFT. Therefore, we propose using the longest CoT data in the SFT phase and the shortest CoT data in the DPO phase.
Specifically, we extract all valid reasoning paths that lead to correct answers from the CoT trees, as multiple correct paths often exist. From these, we select paths categorized by their relative length(short, medium, and long) based on token count, correct paths serve as positive examples, whereas negative examples are generated by identifying incorrect paths that share the shortest common prefix with their corresponding positive paths.
We recognize that "short" and "long" are inherently relative and that CoT lengths vary significantly with problem complexity. For instance, a simple arithmetic problem like "1+1=?" naturally involves a shorter CoT compared to more complex integrals such as "$\int_{0}^{1}\frac{\ln(x+1)}{x^{2}+1},dx=?$". To address these variations systematically, we employ Monte Carlo Tree Search (MCTS) to sample multiple reasoning paths per query and select representative paths by their relative lengths, rather than imposing rigid token-count thresholds.
@@ 3.2 Fine-grained DPO
Llama-3.2-1B-Instruct 43.4 33.0 0.0 0.2 31.9 47.8 33.1
+ Sky-T1 37.2 19.0 0.0 0.0 8.1 10.0 7.5
+ Our Data 50.0 38.6 0.0 5.6 41.1 50.6 47.1
Qwen2.5-7B-Instruct 90.4 62.0 15.0 10.6 69.6 72.8 74.4
+ Sky-T1 89.6 61.6 9.4 0.4 26.2 24.5 30.6
+ Our Data 90.7 64.0 15.0 12.0 73.1 73.4 78.8
Qwen2.5-1.5B-Instruct 67.5 38.4 0.0 1.0 49.6 43.8 41.0
+ Sky-T1 66.7 34.4 0.0 0.0 9.2 11.8 14.8
+ Our Data 74.6 46.8 0.0 5.4 51.5 51.6 54.8
@@ 4.2 Experimental Results
We apply our constructed CoT data to smaller-scale models from the Llama and Qwen families, comparing the results against the Sky-T1 dataset, which employs a QwQ-based distillation pipeline effective on larger models (32B). While Sky-T1 demonstrates competitive performance on large models, it faces challenges when scaled down to 8B models due to the inherent limitations in context processing and reasoning capabilities. In contrast, our CoT data, specifically designed to address these limitations, leads to substantial improvements in smaller models, particularly in tasks involving arithmetic reasoning such as GSM8K and MATH, as well as more complex open-ended tasks like AIME and Blocksworld, as shown in Table 4. This validates the effectiveness of our constructed data in advancing the performance of small models across a wide range of reasoning tasks.
@@ 4.4 MCTS Inference Exploration
As an additional experiment, we explored the impact of applying MCTS at the inference stage. As shown in Table 7, we use Test@N to denote the percentage of problems solved correctly at least once when allowing the model to make N separate guesses for each problem Cobbe et al. (2021b). We evaluated solve rates at Test@1, Test@8, and Test@32 on the MATH dataset. Specifically, at Test@8 and Test@32, the MCTS-based approach outperforms the model without MCTS inference, demonstrating its ability to expand the solution space and leverage test-time scaling effectively.
@@ 5 Related Work
Researchers focus on transferring the reasoning capabilities of LRMs into smaller models, through distillation techniques. Distillation methods, such as fine-tuning smaller models on CoT data generated by larger models Huang et al. (2024a), show that small models can benefit from reasoning data generated by large models. However, recent findings reveal that small models often fail to capture intricate reasoning patterns due to their limited capacity, leading to suboptimal performance when directly distilled from large models Li et al. (2025a); Wang et al. (2025a).
@@ 6 Conclusion
We explore strategies to transfer long CoT reasoning to smaller models, addressing learning difficulty and bias inheritance in distillation. We propose a MCTS framework that independently generates flexible, tree-based reasoning data, reducing reliance on large teacher models. Enhanced by CoT-aware post-training, our approach effectively mitigates overly formalistic long-time thinking. Experiments across diverse tasks (math, planning and instruction-following) demonstrate robust performance improvements. These findings highlight the importance of well-designed data and post-training strategies in improving the efficiency and reliability of smaller-scale reasoning models. In the future, we will continue to explore how reasoning techniques can improve a broader range of non-mathematical tasks, including machine translation Liu et al. (2025), multilingual and multimodal reasoning Zeng et al. (2025); Li et al. (2025b).